OpenStudy (anonymous):

Help, I will medal :D What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2x+4 (blank, blank)

OpenStudy (mathmate):

hints: if y=(3/2)x, and y=-(1/2)x+4 (notice proper parentheses) then you can equate the two y's to give (3/2)x=-(1/2)x+4 and hence solve for x. Substitute x into one of the two original equations to find y.